# From Lipschitz Embedding to Lipschitz Equivalence between Dust-like Self-similar Sets

### v1

@@ 3.1. Localization for rescaling
where $\delta_{n}:=\frac{L}{\Delta_{F}}\cdot\delta^{n}$. The following lemma records several basic facts that will be used frequently later. Recall that $\underline{r}:=\min_{i\in I}r_{i}$ and $\underline{\lambda}:=\min_{j\in J}\lambda_{j}$.
@@
The first statement follows directly from (3.1). For (2), note that
@@
$\operatorname{dist}(F_{\bm{j}},F_{\bm{j}^{\prime}})\geq\lambda_{\bm{j}^{-}}\Delta_{F}>\delta_{n}\Delta_{F}=\frac{L}{\Delta_{F}}\delta^{n}\Delta_{F}=L\delta^{n}.$
@@
Pick the smallest $n$ so that $\delta_{n}\leq\rho$. Since $\bm{j}_{n}(y)\in J_{n}$, $\operatorname{diam}(F_{\bm{j}_{n}(y)})=\lambda_{\bm{j}_{n}(y)}\leq\delta_{n}$ and thus $F_{\bm{j}_{n}(y)}\subset B(y,\rho)$. On the other hand, by Lemma 3.3(1) and the minimality of $n$ (so $\delta_{n}\approx\rho$), there is a constant $C>0$ (independent of $y,n,\varepsilon$) such that
@@
Combining all these observations and noticing that
$\big{(}F\cap B(y,\rho)\big{)}\setminus f(K)\supset\big{(}F_{\mathbf{j}_{n}(y)}\cap B(y,\rho)\big{)}\setminus f(K)=F_{\mathbf{j}_{n}(y)}\setminus f(K),$
we have
$\mathcal{H}^{s}(f(K)\cap B(y,\rho))$ $\leq\mathcal{H}^{s}(F\cap B(y,\rho))-\mathcal{H}^{s}(F_{\bm{j}_{n}(y)}\setminus f(K))$
@@
where the last inequality follows from (3.4). Recalling (3.3), we immediately obtain
@@
To obtain an effective rescaling, we also need to identify a non-condensed local area of the image set $f(K)$.
@@
There is an integer $Q\geq 1$ satisfying the following property. For any $0<\varepsilon<1$, we can find some $n$ and $\bm{j}\in J_{n}$ such that $\frac{\mathcal{H}^{s}(F_{\bm{j}}\cap f(K))}{\mathcal{H}^{s}(F_{\bm{j}})}>1-\varepsilon$ and
@@
By Lemma 3.3(3), the first equality holds trivially since the two sets are identical. It remains to verify the cardinality. Let $0<\varepsilon<1$. By Lemma 3.4, there exists $E\subset f(K)$ with $\mathcal{H}^{s}(E)=\mathcal{H}^{s}(f(K))$ such that for all $y\in E$,
@@
For convenience, below we simply write $\bm{j}_{n_{t}}:=\bm{j}_{n_{t}}(y_{t})$.
Pick any $p\geq 1$ such that
(3.6) $\mathcal{H}^{s}\Big{(}\bigcup_{t=1}^{p}F_{\bm{j}_{n_{t}}}\Big{)}\geq\frac{1}{2}\mathcal{H}^{s}(E)=\frac{1}{2}\mathcal{H}^{s}(f(K)).$
@@
and $\mathcal{M}:=\bigcup_{t=1}^{p}\mathcal{M}_{t}$. Since $F_{\bm{j}_{n_{1}}},\ldots,F_{\bm{j}_{n_{p}}}$ are distinct, $\mathcal{M}_{1},\ldots,\mathcal{M}_{p}$ are disjoint. We claim that $\mathcal{M}\gtrsim\delta_{N}^{-s}$.
@@
(3.9) $=(\underline{\lambda}^{s}-C^{-1}\underline{\lambda}^{-s})\delta_{n_{t}}^{s}\delta_{P}^{-s}=:C_{1}\delta_{n_{t}}^{s}\delta_{P}^{-s}.$
It follows from our choice of $C$ that $C_{1}>0$. Since the above estimate holds for all $1\leq t\leq p$,
$\#\mathcal{M}$ $=\sum_{t=1}^{p}\#\mathcal{M}_{t}$ (by the disjointness)
$\geq C_{1}\delta_{N}^{-s}\sum_{t=1}^{p}\delta_{n_{t}}^{s}$ (by (3.9))
$=C_{1}\delta_{P}^{-s}\cdot\mathcal{H}^{s}(F)^{-1}\sum_{t=1}^{p}\delta_{n_{t}}^{s}\mathcal{H}^{s}(F)$
$\geq C_{1}\delta_{P}^{-s}\cdot\mathcal{H}^{s}(F)^{-1}\sum_{t=1}^{p}\mathcal{H}^{s}(F_{\bm{j}_{n_{t}}})$ (since $\lambda_{\bm{j}_{n_{t}}}\leq\delta_{n_{t}}$)
$\geq C_{1}\delta_{P}^{-s}\cdot\mathcal{H}^{s}(F)^{-1}\cdot\frac{1}{2}\mathcal{H}^{s}(f(K))$ (by (3.6))
$=:C_{2}\delta_{P}^{-s},$
@@
Now, by Lemma 3.3(3), $\{\{\bm{i}\in I_{P}:f(K_{\bm{i}})\cap F_{\eta}\neq\varnothing\}\}_{\eta\in\mathcal{M}}$ is a disjoint family. By the pigeonhole principle, there must exist a word $\eta\in\mathcal{M}$ such that
$\{\bm{i}\in I_{P}:f(K_{\bm{i}})\cap F_{\eta}\neq\varnothing\}\leq\frac{\#I_{P}}{\#\mathcal{M}}.$
@@
$\{\bm{i}\in I_{P}:f(K_{\bm{i}})\cap F_{\eta}\neq\varnothing\}\lesssim\frac{\delta^{-Ps}}{C_{2}\delta_{P}^{-s}}=\frac{\delta^{-Ps}}{C_{2}}\cdot\Big{(}\frac{L\delta^{P}}{\Delta_{F}}\Big{)}^{s}=\frac{L^{s}}{C_{2}\Delta_{F}^{s}},$
@@
Now let us prove Theorem 3.2. Applying Proposition 3.5 to $\varepsilon=n^{-1}$, we find $\{k_{n}\}_{n=1}^{\infty}\subset\mathbb{Z}^{+}$ and $\{\bm{j}_{n}\}_{n=1}^{\infty}$ (say $\bm{j}_{n}\in J_{k_{n}}$) such that
(3.10) $\mathcal{H}^{s}(F_{\bm{j}_{n}}\cap f(K))>(1-n^{-1})\mathcal{H}^{s}(F_{\bm{j}_{n}})$
@@
For convenience, enumerate $\{\bm{i}\in I_{k_{n}}:f(K_{\bm{i}})\subset F_{\bm{j}_{n}}\}=:\{\bm{i}_{n,q}\}_{q=1}^{P}$.
With the notation, for every $n\geq 1$, $f$ sends each $K_{\bm{i}_{n,q}}$ into $F_{\bm{j}_{n}}=\psi_{\bm{j}_{n}}(F)$. This naturally generates a blow-up map
$f_{n}:\bigcup_{q=1}^{P}\psi_{\bm{j}_{n}}^{-1}(K_{\bm{i}_{n,q}})$ $\to F$
@@
It follows that
$\lim_{n\to\infty}\mathcal{H}^{s}\Big{(}F\cap f_{n}\Big{(}\bigcup_{q=1}^{P}\psi_{\bm{j}_{n}}^{-1}(K_{\bm{i}_{n,q}})\Big{)}\Big{)}$ $=\lim_{n\to\infty}\mathcal{H}^{s}\Big{(}F\cap\bigcup_{q=1}^{P}\psi_{\bm{j}_{n}}^{-1}f(K_{\bm{i}_{n,q}})\Big{)}$
$=\mathcal{H}^{s}(F)\cdot\lim_{n\to\infty}\frac{\mathcal{H}^{s}(F\cap\bigcup_{q=1}^{P}\psi_{\bm{j}_{n}}^{-1}f(K_{\bm{i}_{n,q}}))}{\mathcal{H}^{s}(F)}$
$=\mathcal{H}^{s}(F)\cdot\lim_{n\to\infty}\frac{\mathcal{H}^{s}(F_{\bm{j}_{n}}\cap\bigcup_{q=1}^{P}f(K_{\bm{i}_{n,q}}))}{\mathcal{H}^{s}(F_{\bm{j}_{n}})}$
(3.12) $=\mathcal{H}^{s}(F),$
where the last equality follows from (3.11).
For $1\leq q\leq P$, write $\psi_{\bm{j}_{n}}^{-1}(K_{\bm{i}_{n,q}})=:c_{n,q}O_{n,q}K+t_{n,q}$, where $c_{n,q}:=\lambda_{\bm{j}_{n}}^{-1}r_{\bm{i}_{n,q}}$. Recall that $\bm{j}_{n}\in J_{k_{n}}$ and $\bm{i}_{n,q}\in I_{k_{n}}$. By Lemma 3.3, it is easy to compute that $c_{n,q}\in[\frac{\underline{r}\Delta_{F}}{L},\frac{\Delta_{F}}{\underline{\lambda}L}]$ and
$\min_{q\neq q^{\prime}}\operatorname{dist}(\psi_{\bm{j}_{n}}^{-1}(K_{\bm{i}_{n,q}}),\psi_{\bm{j}_{n}}^{-1}(K_{\bm{i}_{n,q^{\prime}}}))$ $=\lambda_{\bm{j}_{n}}^{-1}\min_{q\neq q^{\prime}}\operatorname{dist}(K_{\bm{i}_{n,q}},K_{\bm{i}_{n,q^{\prime}}})$
@@
Applying Lemma 2.2, we can find a Lipschitz map $g:\bigcup_{q=1}^{P}(c_{q}O_{q}K+a_{q})\to F$, where $c_{q}\in[\frac{\underline{r}\Delta_{F}}{L},\frac{\Delta_{F}}{\underline{\lambda}L}]$, $O_{q}$ is a $d\times d$ orthogonal matrix and $a_{q}\in\mathbb{R}^{d}$, such that $\operatorname{Lip}(g)\leq L$ and (recalling (3.12))
$\mathcal{H}^{s}\Big{(}g\Big{(}\bigcup_{q=1}^{P}(c_{q}O_{q}K+a_{q})\Big{)}\Big{)}\geq\lim_{n\to\infty}\mathcal{H}^{s}\Big{(}F\cap f_{n}\Big{(}\bigcup_{q=1}^{P}\psi_{\bm{j}_{n}}^{-1}(K_{\bm{i}_{n,q}})\Big{)}\Big{)}=\mathcal{H}^{s}(F).$
So $g(\bigcup_{q=1}^{P}(c_{q}O_{q}K+a_{q}))$ is a closed subset of $F$ with full measure. As a result, $F=g(\bigcup_{q=1}^{P}(c_{q}O_{q}K+a_{q}))$.
By Baire’s theorem, there is some $q$ such that $g(cO_{q}K+a_{q})$ contains an interior part of $F$, say $F_{\bm{j}}\subset g(cO_{q}K+a_{q})$. Thus, similarly as in the proof of Lemma 3.1, it is easy to construct a Lipschitz map sending $cO_{q}K+a_{q}$ onto $F_{\bm{j}}$. Since $cO_{q}K+a_{q}$ and $F_{\bm{j}}$ are nothing but scaled copies of $K$ and $F$, respectively, this establishes Theorem 3.2.
@@ 4. Lipschitz surjections with measure-preserving properties
In this section, we fix a Lipschitz embedding $f:K\to F$. By Theorem 3.2, we may assume $f$ to be surjective. Again write $L:=\operatorname{Lip}(f)$ for simplicity. A key insight, noted in [3, 21], asserts that any bi-Lipschitz map between $K$ and $F$ (if there is any) should preserve the $s$-dimensional Hausdorff measure in some local area. See [14, Lemma 2.4] for a short proof. While the given map $f$ may fall short of this ideal, we can nevertheless produce an alternative Lipschitz surjection from $K$ to $F$ that enjoys some nice measure linearity.
@@
It remains to check that $\bigcup_{\bm{i}\in\mathcal{I}}\psi_{\bm{j}}^{-1}(K_{\bm{i}})\in\mathcal{K}$. By Lemma 3.3(1), we have
@@
The above lemma tells us that the infimum is well defined. Then one can take $\{E_{n}\}\subset\mathcal{K}$ and a sequence $\{g_{n}\}$, where each $g_{n}:E_{n}\to F$ is a Lipschitz surjection with $\operatorname{Lip}(g_{n})\leq L$, such that $\lim_{n}\mathcal{H}^{s}(E_{n})$ equals the infimum. By Lemma 2.2, we can find $\widetilde{K}\in\mathcal{K}$ and a Lipschitz map $\widetilde{f}:\widetilde{K}\to F$ with $\operatorname{Lip}(\widetilde{f})\leq L$ such that
@@
In the sequel, let us fix such an optimal domain $\widetilde{K}\in\mathcal{K}$, say $\widetilde{K}=\bigcup_{t=1}^{p}(a_{t}O_{t}K+b_{t})$, and a Lipschitz surjection $\widetilde{f}:\widetilde{K}\to F$ with $\operatorname{Lip}(\widetilde{f})\leq L$. Write $\widetilde{c}:=\frac{\mathcal{H}^{s}(\widetilde{K})}{\mathcal{H}^{s}(F)}$.
@@
(4.3) $I^{\prime}_{t,n}:=\{\bm{i}\in I^{*}:a_{t}r_{\bm{i}}\leq\delta^{n}<a_{t}r_{\bm{i}^{-}}\},\quad 1\leq t\leq p.$
@@
For every $n$ and $\bm{j}\in J_{n}$, $\widetilde{f}^{-1}(F_{\bm{j}})=\bigcup_{t=1}^{p}\bigcup_{\bm{i}\in\mathcal{I}_{t}}(a_{t}O_{t}K_{\bm{i}}+b_{t})$ for some $\mathcal{I}_{t}\subset I^{\prime}_{t,n}$. Moreover, if $\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}}))}{\mathcal{H}^{s}(F_{\bm{j}})}\leq\widetilde{c}$, then $\psi_{\bm{j}}^{-1}(\widetilde{f}^{-1}(F_{\bm{j}}))\in\mathcal{K}$ and $\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}}))}{\mathcal{H}^{s}(F_{\bm{j}})}=\widetilde{c}$.
@@
$=La_{t}r_{\bm{i}}\leq L\delta^{n}<\min_{\eta\neq\eta^{\prime}\in J_{n}}\operatorname{dist}(F_{\eta},F_{\eta^{\prime}}),$
@@
(4.4) $\widetilde{f}^{-1}(F_{\bm{j}})=\bigcup_{t=1}^{p}\bigcup_{\bm{i}\in\mathcal{I}_{t}}(a_{t}O_{t}K_{\bm{i}}+b_{t})$
for some $\mathcal{I}_{t}\subset I^{\prime}_{t,n}$.
Secondly, we show that $\psi_{\bm{j}}^{-1}\widetilde{f}^{-1}(F_{\bm{j}})=\bigcup_{t=1}^{m}\bigcup_{\bm{i}\in\mathcal{I}_{t}}\psi_{\bm{j}}^{-1}(a_{t}O_{t}K_{\bm{i}}+b_{t})$ is a member of $\mathcal{K}$ provided that $\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}}))}{\mathcal{H}^{s}(F_{\bm{j}})}\leq\widetilde{c}$. Let $1\leq t,t^{\prime}\leq p$, $\bm{i}\in\mathcal{I}_{t}$ and $\bm{i}^{\prime}\in\mathcal{I}_{t^{\prime}}$. By Lemma 3.3(1), every ratio $\lambda_{\bm{j}}^{-1}a_{t}r_{\bm{i}}\in[\frac{\underline{r}\Delta_{F}}{L},\frac{\Delta_{F}}{\underline{\lambda}L}]$. On the other hand, if $t\neq t^{\prime}$ then since $\widetilde{K}\in\mathcal{K}$,
@@
$>\operatorname{dist}(a_{t}O_{t}K+b_{t},a_{t^{\prime}}O_{t^{\prime}}K+b_{t^{\prime}})\geq\gamma=L^{-1}\Delta_{F}\Delta_{K};$

### v2

@@ 3.1. Localization for rescaling
where $\delta_{n}:=\frac{L}{\Delta_{F}}\cdot\delta^{n}$ is a constant multiple of $\delta^{n}$. The next lemma records several basic facts that will be used frequently later. Recall that $\underline{r}:=\min_{i\in I}r_{i}$ and $\underline{\lambda}:=\min_{j\in J}\lambda_{j}$.
@@
The first statement follows directly from (3.1) and the fact that $\sum_{\bm{i}\in I_{n}}r_{\bm{i}}^{s}=\sum_{\bm{j}\in J_{n}}\lambda_{\bm{j}}^{s}=1$. For (2), note that
@@
$\operatorname{dist}(F_{\bm{j}},F_{\bm{j}^{\prime}})\geq\lambda_{\bm{j}^{-}}\Delta_{F}>\delta_{n}\Delta_{F}=L\delta^{n}.$
@@
Pick the smallest $n$ so that $\delta_{n}\leq\rho$. Since $\bm{j}_{n}(y)\in J_{n}$, $\operatorname{diam}(F_{\bm{j}_{n}(y)})=\lambda_{\bm{j}_{n}(y)}\leq\delta_{n}$ and thus $F_{\bm{j}_{n}(y)}\subset B(y,\rho)$. On the other hand, by Lemma 3.3(1) and the minimality of $n$, we have $\rho/C_{1}\leq\delta_{n}\leq C_{1}\rho$ for some constant $C_{1}>1$. By Lemma 3.3(2), one can find a constant $C_{2}>0$ such that for all $n\geq 1$ and $\bm{j}\in J_{n}$, there exists $x_{\bm{j}}\in F_{\bm{j}}$ such that $\{B(x_{\bm{j}},C_{2}\rho)\}_{\bm{j}\in J_{n}}$ are mutually disjoint. Thus
$\#\big{\{}\bm{j}\in J_{n}:F_{\bm{j}}\cap B(y,\rho)\not$$\emptyset\big{\}}$ $\leq\#\big{\{}\bm{j}\in J_{n}:B(x_{\bm{j}},C_{2}\rho)\subset B\big{(}y,(1+C_{1}+C_{2})\rho\big{)}\big{\}}$
$\leq\Big{(}\frac{1+C_{1}+C_{2}}{C_{2}}\Big{)}^{d}.$
As a result, there is a constant $C>0$ (independent of $y,n,\varepsilon$) such that
@@
Recalling that $f(K)\subset F$ and $F_{\bm{j}_{n}(y)}\subset B(y,\rho)\cap F$, we have
$\mathcal{H}^{s}(f(K)\cap B(y,\rho))$ $\leq\mathcal{H}^{s}(F\cap B(y,\rho))-\mathcal{H}^{s}((F\cap B(y,\rho))\setminus f(K))$
$\leq\mathcal{H}^{s}(F\cap B(y,\rho))-\mathcal{H}^{s}(F_{\bm{j}_{n}(y)}\setminus f(K))$
@@
where the last inequality follows from (3.4). By (3.3), we immediately obtain
@@
To obtain an effective rescaling, we also need to determine a local region within $F$ where the map $f$ is not overly concentrated.
@@
There is an integer $Q\geq 1$ satisfying the following property. For any $0<\varepsilon<1$, we can find some $n$ and $\bm{j}\in J_{n}$ such that $\frac{\mathcal{H}^{s}(f(K)\cap F_{\bm{j}})}{\mathcal{H}^{s}(F_{\bm{j}})}>1-\varepsilon$ and
@@
The first equality is immediately from Lemma 3.3(3) as the two sets are identical. It remains to verify the cardinality. Let $0<\varepsilon<1$. By Lemma 3.4, there exists $E\subset f(K)$ with $\mathcal{H}^{s}(E)=\mathcal{H}^{s}(f(K))$ such that for all $y\in E$,
@@
For convenience, below we abbreviate $\bm{j}_{n_{t}}:=\bm{j}_{n_{t}}(y_{t})$.
Pick any $p\geq 1$ such that $\mathcal{H}^{s}(\bigcup_{t=1}^{p}F_{\bm{j}_{n_{t}}})\geq\mathcal{H}^{s}(E)/2=\mathcal{H}^{s}(f(K))/2$. So
(3.6) $\sum_{t=1}^{p}\delta_{n_{t}}^{s}\geq\mathcal{H}^{s}(F)^{-1}\sum_{t=1}^{p}\lambda_{n_{t}}^{s}\mathcal{H}^{s}(F)=\mathcal{H}^{s}(F)^{-1}\sum_{t=1}^{p}\mathcal{H}^{s}(F_{\bm{j}_{n_{t}}})\geq\frac{\mathcal{H}^{s}(f(K))}{2\mathcal{H}^{s}(F)}.$
@@
and $\mathcal{M}:=\bigcup_{t=1}^{p}\mathcal{M}_{t}$. Since $F_{\bm{j}_{n_{1}}},\ldots,F_{\bm{j}_{n_{p}}}$ are disjoint, $\mathcal{M}_{1},\ldots,\mathcal{M}_{p}$ are disjoint. We claim that $\mathcal{M}\gtrsim\delta_{P}^{-s}$.
@@
(3.9) $=(\underline{\lambda}^{s}-C^{-1}\underline{\lambda}^{-s})\delta_{n_{t}}^{s}\delta_{P}^{-s}=:C_{1}\delta_{n_{t}}^{s}\delta_{P}^{-s},$
where $C_{1}>0$ since $C>\underline{\lambda}^{-2s}$. Letting $C_{2}:=\frac{C_{1}\mathcal{H}^{s}(f(K))}{2\mathcal{H}^{s}(F)}$, we know from the disjointness of $\{\mathcal{M}_{t}\}_{t=1}^{p}$, the estimates (3.9) and (3.6) that
$\#\mathcal{M}=\sum_{t=1}^{p}\#\mathcal{M}_{t}\geq C_{1}\delta_{P}^{-s}\sum_{t=1}^{p}\delta_{n_{t}}^{s}\geq C_{2}\delta_{P}^{-s},$
@@
Lemma 3.3(3) also implies that $\{\{\bm{i}\in I_{P}:f(K_{\bm{i}})\cap F_{\eta}\neq\varnothing\}\}_{\eta\in\mathcal{M}}$ is a disjoint family. By the pigeonhole principle, there must exist a word $\eta_{0}\in\mathcal{M}$ such that
$\{\bm{i}\in I_{P}:f(K_{\bm{i}})\cap F_{\eta_{0}}\neq\varnothing\}\leq\frac{\#I_{P}}{\#\mathcal{M}}.$
@@
$\{\bm{i}\in I_{P}:f(K_{\bm{i}})\cap F_{\eta_{0}}\neq\varnothing\}\lesssim\frac{\delta^{-Ps}}{C_{2}\delta_{P}^{-s}}=\frac{\delta^{-Ps}}{C_{2}}\cdot\Big{(}\frac{L\delta^{P}}{\Delta_{F}}\Big{)}^{s}=\frac{L^{s}}{C_{2}\Delta_{F}^{s}},$
@@
Now let us prove Theorem 3.2. Applying Proposition 3.5 to $\varepsilon=n^{-1}$, we find for all $n\geq 1$ an integer $k_{n}\geq 1$ and a word $\bm{j}_{n}\in J_{k_{n}}$ such that
(3.10) $\mathcal{H}^{s}(F_{\bm{j}_{n}}\cap f(K))>(1-n^{-1})\mathcal{H}^{s}(F_{\bm{j}_{n}}),$
@@
For convenience, enumerate $\{\bm{i}\in I_{k_{n}}:f(K_{\bm{i}})\subset F_{\bm{j}_{n}}\}=:\{\bm{i}_{n,t}\}_{t=1}^{P}$.
With the notation, for every $n\geq 1$, $f$ sends each $K_{\bm{i}_{n,t}}$ into $F_{\bm{j}_{n}}=\psi_{\bm{j}_{n}}(F)$. This naturally generates a blow-up map
$f_{n}:\bigcup_{t=1}^{P}\psi_{\bm{j}_{n}}^{-1}(K_{\bm{i}_{n,t}})$ $\to F$
@@
Clearly, $\operatorname{Lip}(f_{n})\leq\operatorname{Lip}(f)=L$ for all $n$. By (3.11),
$\lim_{n\to\infty}\frac{\mathcal{H}^{s}\big{(}F\cap f_{n}(\bigcup_{t=1}^{P}\psi_{\bm{j}_{n}}^{-1}(K_{\bm{i}_{n,t}}))\big{)}}{\mathcal{H}^{s}(F)}$ $=\lim_{n\to\infty}\frac{\mathcal{H}^{s}\big{(}F\cap\bigcup_{t=1}^{P}\psi_{\bm{j}_{n}}^{-1}f(K_{\bm{i}_{n,t}})\big{)}}{\mathcal{H}^{s}(F)}$
(3.12) $=\lim_{n\to\infty}\frac{\mathcal{H}^{s}\big{(}F_{\bm{j}_{n}}\cap\bigcup_{t=1}^{P}f(K_{\bm{i}_{n,t}})\big{)}}{\mathcal{H}^{s}(F_{\bm{j}_{n}})}=1.$
For $1\leq t\leq P$, write $\psi_{\bm{j}_{n}}^{-1}(K_{\bm{i}_{n,t}})=:c_{n,t}O_{n,t}K+b_{n,t}$, where $c_{n,t}:=\lambda_{\bm{j}_{n}}^{-1}r_{\bm{i}_{n,t}}$. Recall that $\bm{j}_{n}\in J_{k_{n}}$ and $\bm{i}_{n,t}\in I_{k_{n}}$. By Lemma 3.3, it is easy to verify that $c_{n,t}\in[\frac{\underline{r}\Delta_{F}}{L},\frac{\Delta_{F}}{\underline{\lambda}L}]$ and
$\min_{t\neq t^{\prime}}\operatorname{dist}(\psi_{\bm{j}_{n}}^{-1}(K_{\bm{i}_{n,t}}),\psi_{\bm{j}_{n}}^{-1}(K_{\bm{i}_{n,t^{\prime}}}))$ $=\lambda_{\bm{j}_{n}}^{-1}\min_{t\neq t^{\prime}}\operatorname{dist}(K_{\bm{i}_{n,t}},K_{\bm{i}_{n,t^{\prime}}})$
@@
Applying Lemma 2.2, we can find a Lipschitz map $g:\bigcup_{t=1}^{P}(c_{t}O_{t}K+b_{t})\to F$, where $c_{t}\in[\frac{\underline{r}\Delta_{F}}{L},\frac{\Delta_{F}}{\underline{\lambda}L}]$, $O_{t}$ is a $d\times d$ orthogonal matrix and $b_{t}\in\mathbb{R}^{d}$, such that $\operatorname{Lip}(g)\leq L$ and (recalling (3.12))
$\mathcal{H}^{s}\Big{(}g\Big{(}\bigcup_{t=1}^{P}(c_{t}O_{t}K+b_{t})\Big{)}\Big{)}\geq\lim_{n\to\infty}\mathcal{H}^{s}\Big{(}F\cap f_{n}\Big{(}\bigcup_{t=1}^{P}\psi_{\bm{j}_{n}}^{-1}(K_{\bm{i}_{n,t}})\Big{)}\Big{)}=\mathcal{H}^{s}(F).$
Therefore, the image $g(\bigcup_{t=1}^{P}(c_{t}O_{t}K+b_{t}))$ is a closed subset of $F$ with full measure. It follows that $g(\bigcup_{t=1}^{P}(c_{t}O_{t}K+b_{t}))=F$.
By Baire’s theorem, there is some $t$ such that $g(c_{t}O_{t}K+b_{t})$ contains an interior part of $F$, say $F_{\bm{j}}\subset g(c_{t}O_{t}K+b_{t})$. Then, similarly as in the proof of Lemma 3.1, it is easy to construct a Lipschitz map sending $c_{t}O_{t}K+b_{t}$ onto $F_{\bm{j}}$. Since $c_{t}O_{t}K+b_{t}$ and $F_{\bm{j}}$ are nothing but scaled copies of $K$ and $F$, respectively, this establishes Theorem 3.2.
@@ 4. Lipschitz surjections with measure-preserving properties
In this section, we fix a Lipschitz embedding $f:K\to F$ with $\mathcal{H}^{s}(f(K))>0$. By Theorem 3.2, we may assume $f$ to be surjective. Again write $L:=\operatorname{Lip}(f)$ for simplicity. A key insight, noted in [3, 21], asserts that any bi-Lipschitz map between $K$ and $F$ (if there is any) should preserve the $s$-dimensional Hausdorff measure in some local area; see [14, Lemma 2.4] for a short proof. While the given map $f$ may fall short of this ideal, we can nevertheless produce an alternative Lipschitz surjection from $K$ to $F$ that enjoys some nice measure linearity.
@@
It remains to check that $\bigcup_{\bm{i}\in\mathcal{I}}\psi_{\bm{j}}^{-1}(K_{\bm{i}})\in\mathcal{K}$. By Lemma 3.3, we have
@@
The above lemma tells us that the infimum is well defined. Then one can take $\{E_{n}\}_{n=1}^{\infty}\subset\mathcal{K}$ and a sequence $\{g_{n}\}_{n=1}^{\infty}$, where each $g_{n}:E_{n}\to F$ is a Lipschitz surjection with $\operatorname{Lip}(g_{n})\leq L$, such that $\lim_{n}\mathcal{H}^{s}(E_{n})$ equals the infimum. By Lemma 2.2, we can find $\widetilde{K}\in\mathcal{K}$ and a Lipschitz map $\widetilde{f}:\widetilde{K}\to F$ with $\operatorname{Lip}(\widetilde{f})\leq L$ such that
@@
In the sequel of the paper, let us fix such an optimal $\widetilde{K}\in\mathcal{K}$, say $\widetilde{K}=\bigcup_{t=1}^{p}(a_{t}O_{t}K+b_{t})$, and a Lipschitz surjection $\widetilde{f}:\widetilde{K}\to F$ with $\operatorname{Lip}(\widetilde{f})\leq L$. Write $\widetilde{c}:=\frac{\mathcal{H}^{s}(\widetilde{K})}{\mathcal{H}^{s}(F)}$. Also fix a small constant $0<\delta<\min\{1,a_{1},a_{2},\ldots,a_{p}\}$.
@@
(4.3) $I^{\prime}_{t,n}:=\{\bm{i}\in I^{*}:a_{t}r_{\bm{i}}\leq\delta^{n}<a_{t}r_{\bm{i}^{-}}\},\quad 1\leq t\leq p$
@@
For every $n$ and $\bm{j}\in J_{n}$, $\widetilde{f}^{-1}(F_{\bm{j}})=\bigcup_{t=1}^{p}\bigcup_{\bm{i}\in\mathcal{I}_{t}}(a_{t}O_{t}K_{\bm{i}}+b_{t})$ with $\mathcal{I}_{t}\subset I^{\prime}_{t,n}$ for all $1\leq t\leq p$. Moreover, if $\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}}))}{\mathcal{H}^{s}(F_{\bm{j}})}\leq\widetilde{c}$, then $\psi_{\bm{j}}^{-1}(\widetilde{f}^{-1}(F_{\bm{j}}))\in\mathcal{K}$ and $\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}}))}{\mathcal{H}^{s}(F_{\bm{j}})}=\widetilde{c}$.
@@
(4.4) $=La_{t}r_{\bm{i}}\leq L\delta^{n}<\min_{\eta\neq\eta^{\prime}\in J_{n}}\operatorname{dist}(F_{\eta},F_{\eta^{\prime}}),$
@@
(4.5) $\widetilde{f}^{-1}(F_{\bm{j}})=\bigcup_{t=1}^{p}\bigcup_{\bm{i}\in\mathcal{I}_{t}}(a_{t}O_{t}K_{\bm{i}}+b_{t}),$
where $\mathcal{I}_{t}\subset I^{\prime}_{t,n}$ for all $1\leq t\leq p$.
Secondly, we show that $\psi_{\bm{j}}^{-1}\widetilde{f}^{-1}(F_{\bm{j}})=\bigcup_{t=1}^{p}\bigcup_{\bm{i}\in\mathcal{I}_{t}}\psi_{\bm{j}}^{-1}(a_{t}O_{t}K_{\bm{i}}+b_{t})$ is a member of $\mathcal{K}$ provided that $\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}}))}{\mathcal{H}^{s}(F_{\bm{j}})}\leq\widetilde{c}$. Let $1\leq t,t^{\prime}\leq p$, $\bm{i}\in\mathcal{I}_{t}$ and $\bm{i}^{\prime}\in\mathcal{I}_{t^{\prime}}$. By Lemma 3.3(1), every ratio $\lambda_{\bm{j}}^{-1}a_{t}r_{\bm{i}}\in[\frac{\underline{r}\Delta_{F}}{L},\frac{\Delta_{F}}{\underline{\lambda}L}]$. Furthermore, if $t\neq t^{\prime}$ then since $\widetilde{K}\in\mathcal{K}$,
@@
$>\operatorname{dist}(a_{t}O_{t}K_{\bm{i}}+b_{t},a_{t^{\prime}}O_{t^{\prime}}K_{\bm{i}^{\prime}}+b_{t^{\prime}})$
$\geq\operatorname{dist}(a_{t}O_{t}K+b_{t},a_{t^{\prime}}O_{t^{\prime}}K+b_{t^{\prime}})\geq\gamma;$